# Enhancing Content-based Recommendation via Large Language Model

### v1

@@ 3.5. Model Optimization
Finally, we concatenate $\mathbf{v}_{u}$ and $\mathbf{v}_{i}$ to obtain the final score:
(9) $\small\begin{split}\hat{r}_{\mathrm{u,i}}=\text{Predict}(\text{Concat}(\mathbf{v}_{u},\mathbf{v}_{i})),\ \ \mathcal{L}_{\mathrm{rec}}=\frac{1}{{|\mathcal{D}_{\mathrm{target}}|}}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{target}}}(\hat{r}_{\mathrm{u,i}}-r_{\mathrm{u,i}})^{2},\end{split}$
@@
Throughout the training process, the total loss comprises the MSE loss of rating prediction loss and the contrastive loss incurred during contrastive learning:
(10) $\small\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\lambda\mathcal{L}_{\mathrm{cl}},$
@@ 4.1. Experimental Setup
Datasets DeepConn D-attn ALFM ANR BiGI RGCL LoID LoID(Elec) LoID(Elec &Movie)
Electronics 1.659 1.744 1.563 1.445 1.1433 0.9621 0.8662 $\star$ $\star$
Movies & TV 1.207 1.246 1.193 1.112 1.3489 0.9239 0.8353 0.8167 $\star$
CDs & Vinyl 0.980 1.014 0.956 0.914 1.1831 0.8180 0.6516 0.6295 0.6217
Amazon Instant Video 1.178 1.213 1.075 1.009 1.2437 0.9357 0.5177 0.4891 0.4864
Baby 1.442 1.507 1.359 1.258 1.5166 1.1414 0.7373 0.6911 0.6876
Digital Music 0.749 0.775 0.725 0.688 0.7731 0.7735 0.4696 0.4380 0.4365
Musical Instruments 1.160 1.224 1.072 1.034 1.3580 0.7211 0.5830 0.5179 0.5166
Office Products 1.569 1.650 1.474 1.337 1.6902 0.7001 0.6150 0.5909 0.5573
Patio, Lawn & Garden 1.622 1.696 1.510 1.403 1.7205 0.7049 0.6622 0.6358 0.6129
Pet Supplies 1.565 1.628 1.485 1.377 1.6185 1.2380 0.7685 0.7304 0.7249
Video Games 1.498 1.533 1.383 1.292 1.4450 1.0826 0.7822 0.7541 0.7409
Remark $\star$’ indicates that this domain is served as the source domain.
@@ 4.1.2. Baselines
We compare LoID with several baselines, which can be categorized into three classes: (1) Single-aspect methods, such as DeepCoNN (Zheng et al., 2017) and D-Attn (Seo et al., 2017), extract features from the historical semantic information of users and items using CNN and attention mechanisms. (2) Multi-aspect methods, including ALFM (Cheng et al., 2018) and ANR (Chin et al., 2018), aim to extract multiple semantic aspects and their respective importance. (3) GNNs-based methods, such as BiGI (Cao et al., 2021) and RGCL (Shuai et al., 2022), integrate adjacent node information from collaborative filtering.
4.1.3. Parameter Settings
In our method, the embedding size $d$ is fixed as 768, the dropout rate is fixed 0.5, the learning rate is set as 1e-5, the batch size is fixed as 4, and the training spans 5 epochs. In the LoRA module, the low-rank hyper-parameter $r$ selected from 16 to 48 with step length 8; the $\lambda$ selected from 0.2 to 0.5 with step length 0.1; the number of extracted user/item historical contents $k$ is is set as $[3\rightarrow 5\rightarrow 7\rightarrow 10]$. In the following section, we report LoID results under $r=32$, $\lambda=0.2$, and $k=5$ by default.
In the experiment, we report the DeepConn, D-attn, ALFM, and ANR results in the original literature (Chin et al., 2018) directly. Besides, in BiGI, the feature and hidden layer dimension are both set to 128, the dropout rate is fixed to 0.3, the step size is 0.001 and the number of GNN layers is set to 2. In RGCL, the number of GCN aggregation units and the number of output units of RGCL are both set to 64. The dropout rates of Edges, GCN, and nodes are set to 1.0, 0.7, and 0.3. Among our method and all baselines, we use the Adam (Kingma and Ba, 2015) algorithm to update parameters.
@@ 4.2. Performance Comparisons
Table 1 shows the performance of LoID on eleven datasets in terms of MSE metrics. Compared to other baselines, we note that aspect-aware recommendation methods such as ALFM and ANR consistently outperform DeepCoNN and D-Attn. We attribute this to the limitations of DeepCoNN and D-Attn, which lack a comprehensive model for the intricate decision-making process in user-item interactions. In addition, RGCL outperforms single-aspect and multi-aspect methods, indicating superior performance of graph neural networks. However, BiGI, despite incorporating graph networks, falls short compared to single-aspect methods. This suggests incorporating user textual semantic information is effective. We observe that LoID achieves statistically significant improvement over all SOTA baseline methods. This highlights that when using user contents, ID information should not be overlooked; instead, aligning them with semantic information in the same space can achieve positive performance. Moreover, LoID (Elec) results are superior to LoID, emphasizing enhanced performance with information from different domains. This demonstrates the effectiveness of extracting multi-aspect semantic information to empower various domains and our plugin idea effectiveness.
Multi-LoRA Merging. When we employ multiple LoRAs from different domains, the results of Multi-LoRA are superior to Single LoRA. We believe the reason is that multiple LoRAs bring richer and more extensive users’ data, more accurately describing historical behavior and preference tendencies.
@@ 4.3. Discussion of Domain Correlation Effect
To explore the semantic relationship of different domains, we conducted a more granular study of how to enhance further the transfer of semantic information to empower different domains. We consider whether the recommendation performance is correlated to the similarity between the target and source domains. First, we select 100 reviews from each dataset randomly, employ the Sentence-BERT (Reimers and Gurevych, 2019), and quantify the cosine similarity between datasets. Then, we assign different datasets as source domains to show the improvements compared with the origin LoID (as shown in Table 2). We can conclude that an increase in the similarity between domains leads to a corresponding rise in the performance improvement ratio. This implies that the transfer of our method becomes more effective when domains exhibit higher similarity.
@@ 4.4. Discussion of Different LLMs
This section investigates the impact of different LLMs in the pre-training phase, considering two distinct paradigms, i.e., GPT and BERT. Due to prompt-based GPT being unable to predict floating-point numbers, thus we formulate the prompt:
Input template: Give some example: {content1} is score {score1}, {content2} is score {score2}. Guess the score (The score should be between 1 and 5, where 1 means the lowest score, and 5 means the highest score) of {current content}, we think the score is?     Target template: {score}, {explanation}.
To ensure a fair comparison, we tuned them by LoRA and adopted PRF (Precision, Recall, and F1-score) as the evaluation protocol. Table 3 presents the results, indicating that the BERT outperforms GPT2-medium. This superiority may be attributed to BERT belonging to bi-directional encoder-framework LLM, which is more powerful under the regression task than decoder-framework LLM.

### v2

@@ 3.5. Model Optimization
Finally, we concatenate $\mathbf{v}_{u}$ and $\mathbf{v}_{i}$ to predict the rating:
(10) $\footnotesize\begin{split}\hat{r}_{\mathrm{u,i}}=\text{Predict}(\text{Concat}(\mathbf{v}_{u},\mathbf{v}_{i})),\ \ \mathcal{L}_{\mathrm{rec}}=\frac{1}{{|\mathcal{D}_{\mathrm{target}}|}}\sum_{\mathrm{(u,i)}}^{\mathcal{D}_{\mathrm{target}}}(\hat{r}_{\mathrm{u,i}}-r_{\mathrm{u,i}})^{2},\end{split}$
@@
Throughout the training process, the total loss comprises the rating prediction loss and the contrastive loss:
(11) $\footnotesize\mathcal{L}=\mathcal{L}_{\mathrm{rec}}+\lambda\mathcal{L}_{\mathrm{cl}},$
@@ 4.1. Experimental Setup
Datasets DeepConn D-attn ALFM ANR BiGI RGCL LoID(w/o ID CL) LoID LoID(Elec) LoID(Elec &Movie)
Electronics 1.659 1.744 1.563 1.445 1.1433 0.9621 0.8738 0.8666 $\star$ $\star$
Movies & TV 1.207 1.246 1.193 1.112 1.3489 0.9239 0.8526 0.8362 0.8167 $\star$
CDs & Vinyl 0.980 1.014 0.956 0.914 1.1831 0.8180 0.7001 0.6558 0.6295 0.6170
Amazon Instant Video 1.178 1.213 1.075 1.009 1.2437 0.9357 0.5097 0.5094 0.4891 0.4731
Baby 1.442 1.507 1.359 1.258 1.5166 1.1414 0.7436 0.7275 0.6911 0.6888
Digital Music 0.749 0.775 0.725 0.688 0.7731 0.7735 0.4664 0.4709 0.4380 0.4195
Musical Instruments 1.160 1.224 1.072 1.034 1.3580 0.7211 0.5878 0.5816 0.5179 0.5322
Office Products 1.569 1.650 1.474 1.337 1.6902 0.7001 0.6422 0.6016 0.5909 0.5621
Patio, Lawn & Garden 1.622 1.696 1.510 1.403 1.7205 0.7049 0.6833 0.6703 0.6358 0.6281
Pet Supplies 1.565 1.628 1.485 1.377 1.6185 1.2380 0.7436 0.7679 0.7304 0.7196
Video Games 1.498 1.533 1.383 1.292 1.4450 1.0826 0.7994 0.7881 0.7541 0.7336
Remark $\star$’ indicates that this domain is served as the source domain. We report results for LoID and its variants under an average of 5 different random seed result settings.
@@ 4.1.2. Baselines
We compare LoID with three classes of baselines: (1) Single-aspect methods, such as DeepCoNN (Zheng et al., 2017) and D-Attn (Seo et al., 2017), extract features from the historical semantic information of users and items using CNN and attention mechanisms. (2) Multi-aspect methods, including ALFM (Cheng et al., 2018) and ANR (Chin et al., 2018), aim to extract multiple semantic aspects and their respective importance. (3) GNNs-based methods, such as BiGI (Cao et al., 2021) and RGCL (Shuai et al., 2022), integrate adjacent node information from collaborative filtering.
@@
4.1.3. Parameter Settings
The embedding size $d$ ,dropout rate,learning rate and batch size is fixed as 768, 0.5, 1e-5 and 4. In the LoRA module, the low-rank hyper-parameter $r$ selected from 16 to 48 with step length 8; the $\lambda$ selected from 0.2 to 0.5 with step length 0.1; the number of extracted user/item historical contents $k$ is is set as $[3\rightarrow 5\rightarrow 7\rightarrow 10]$. Among all methods, we use the Adam (Kingma and Ba, 2015) to update parameters.
@@ 4.2. Performance Comparisons
Table 1 shows the performance of LoID on eleven datasets. All the experiments are repeated 5 times to sample historical contents in Section 3.4, and we report the (average) test MSE obtained when the validation MSE is the lowest. We observe that LoID achieves significant improvement over all SOTA baseline methods. We note that aspect-aware methods such as ALFM and ANR consistently outperform DeepCoNN and D-Attn. We attribute this to the fact that DeepCoNN and D-Attn lack a comprehensive model for the intricate decision-making process in user-item interactions. In addition, RGCL outperforms single-aspect and multi-aspect methods, indicating superior performance of graph neural networks. However, BiGI, though GNNs-based methods, falls short compared to single-aspect methods, without making use of reviews.
We observe that the results become worse when removing the contrastive learning module. Moreover, LoID (Elec) results are superior to LoID. This demonstrates the effectiveness of our LoRA-based plugin idea. When multiple LoRAs are employed, the results outperform when using single LoRA. We believe more LoRAs, more users’ behavior and preference data.
@@ 4.3. Discussion of Domain Correlation Effect
We investigate whether the performance is correlated to the similarity between the target and source domains. We select $n$ reviews from each dataset randomly, employ the Sentence-BERT (Reimers and Gurevych, 2019), to quantify the cosine similarity between the domains. In our experiments, we set $n$ to be 100. Then, we show the improvements compared with the original LoID (as shown in Table 2). We conclude that as the similarity between the source and target domains increases, there is a corresponding rise in the model’s performance.
@@ 4.4. Discussion of Different LLMs
We investigate the impact of different LLMs in the pre-training phase. We consider two distinct paradigms, i.e., GPT and BERT. Due to prompt-based GPT being unable to predict floating-point numbers, thus we formulate the prompt: Input template: Give some example: {content1} is score {score1}, {content2} is score {score2}. Guess the score (The score should be between 1 and 5, where 1 means the lowest score, and 5 means the highest score) of {current content}, we think the score is?     Target template: {score}, {explanation}. To ensure a fair comparison, we tuned them by LoRA and adopted PRF (Precision, Recall, and F1-score) as the evaluation protocol. Table 3 presents the results, indicating that the BERT outperforms GPT2-medium.
@@ 5. Conclusions
Acknowledgement
This work is supported by the National Natural Science Foundation of China under Grant No. 62206128 and the Undergraduate Research Training Program of Nanjing University of Science and Technology (established in 2023) under Grant No. 2023066021B.